Help Please!
On any given day the probability that a certain gas station will run out of gas is 1/10.
a) What is the probability the gas station will run out of gas one day in the next two weeks?
b) What is the probability the gas station will run out of gas at least two days in the next two weeks?
c) What is the expected number of times the gas station will run out of gas in the next two weeks?

Answer by stanbon(75887)

On any given day the probability that a certain gas station will run out of gas is 1/10.
a) What is the probability the gas station will run out of gas one day in the next two weeks?
n = 14; p = 1/10
P(x=1) = 14C1(1/10)(9/10)^13 = 0.3559
----------------------------------------
b) What is the probability the gas station will run out of gas at least two days in the next two weeks?
P(2<= x <=14) = 1 - P(0<=x<=1) = 1 - 0.4154
----------------------------------------
c) What is the expected number of times the gas station will run out of gas in the next two weeks?
E(x) = np = 14*1/10 = 1.4 times
===================================
Cheers,
Stan H.
==========